OpenStudy (anonymous):

An electric current, I, in amps, is given by I=cos(wt)+8‾‾√sin(wt), where w≠0 is a constant. What are the maximum and minimum values of I?

OpenStudy (anonymous):

I'm not sure how to interpret the current function... Is it \[I(t)=\cos\omega t+8\sqrt{\sin\omega t}\,?\]

OpenStudy (anonymous):

the square root is over the 8 not sin(wt)

OpenStudy (anonymous):

So \(I(t)=\cos\omega t+\sqrt 8\sin\omega t\) ? You can find the maximum and minimum values by taking the derivative and solving for \(I'(t)=0\) (i.e. find the critical points).

OpenStudy (anonymous):

what would the answer be then?

OpenStudy (anonymous):

I'm not here to do your work for you. Here's a starting point: \[I'(t)=-\omega \sin\omega t+\sqrt8\omega\cos\omega t=0~~\iff~~\sqrt8=\tan\omega t\] Solve for \(t\).

OpenStudy (anonymous):

no